OpenStudy (anonymous):

Can somebody help me solve the following? Graph the linear equation y= -1/2x - 5

OpenStudy (jackellyn):

-1/2 is the slope rise -1 --- = ---- run 2 and the y intercept is -5 according to y =mx+b where m is slope and b is y intercept. and to find the x intercept set y =0 and solve for x 0=-1/2x -5 5=-1/2x 5(2)=-x -10=x

OpenStudy (anonymous):

Wow! Thank you so much for the detailed explanation!

OpenStudy (jackellyn):

no problem :)